OpenStudy (anonymous):

What is the area of a circle if its circumference is 20π?

hartnn (hartnn):

Circumference = 2 pi r = 20 pi find r from here then area = pi*r*r

OpenStudy (anonymous):

@hartnn im confuse whats pi

hartnn (hartnn):

pi is an irrational number (approximately 3.14 or 22/7) denoted by \(\huge \pi\)

OpenStudy (anonymous):

oh ok because i got 3.14*2=6.28

hartnn (hartnn):

why multiply by 2??

OpenStudy (mayankdevnani):

20pi = 2pi(r) r = 20pi / 2pi r = 10 units A = pi(10^2) A = 100pi units^2

OpenStudy (mayankdevnani):

@janett77

OpenStudy (anonymous):

@hartnn wait idk do i have to multiply it by 2o

hartnn (hartnn):

i would first like u to solve for r in 2 pi r = 20 pi

OpenStudy (anonymous):

r=10

hartnn (hartnn):

good, now area = pi*r*r= ??

OpenStudy (anonymous):

3.14*10

OpenStudy (anonymous):

is that good so far

hartnn (hartnn):

nopes its 3.14*10*10= 314 approximately.